OpenStudy (anonymous):

Find the product of (x + 5)(x − 5). x^2 − 10x + 25 x^2 + 10x + 25 x^2 − 25 x^2 + 25

OpenStudy (anonymous):

there is a general formula that you must remember \[a ^{2}-b ^{2}=(a-b)(a+b)\]

OpenStudy (anonymous):

Okay, can you tell me what the answer is?

OpenStudy (anonymous):

if u look at the formular it's pretty obvious. in your case a = x and b = 5 (x-5)(x+5) = a^2 - b^2 what are a^2 and b^2? x^2-5^2 = x^2 -25